Consider the following time series data 4 10 13 15 a. Which of the following is a correct time series plot for this data? TimeSeries Value 14 12 10 4 2 TimePeriod(t) TimeSeries Value 14 12 10 4 2 TimePeriod(t) TimeSeries Value 14 7つTimeSeries Value 14 12 10 8 TimePeriod(t) 3. TimeSeries Value 14 12 10 TimePeriod (t -Select your answer+ What type of pattern exists in the data? -Select your answer- b. Develop the linear trend equation for this time series (to 1 decimal). c. What is the forecast for t-6 (to 1 decimal)?
Answer #1

graph 1 is correct

for pattern: please provide option given (it is trend)

b)

(x-X)(y-y) 8.4000 0.2000 0.0000 1.8000 7.6000 18.0000 (x-x) (y-y) 17.6400 0.0400 1.4400 3.2400 14.4400 36.8000 S.No X-X 4.000

trend equation: Tt =5.8+1.8*t

c)

forecast =5.8+1.8*6=16.6